# VRBench: A Benchmark for Multi-Step Reasoning in Long Narrative Videos

### v1

@@ 5 Conclusion
This paper introduces VRBench, a comprehensive long-narrative video benchmark for evaluating multi-step reasoning. Through manually filtered narrative videos, high-quality stepwise annotations, and a multi-phase evaluation pipeline, VRBench distinguishes itself from other existing reasoning benchmarks and shows the robust capability to evaluate LLMs and VLMs both from the process and outcome perspectives. By evaluating and analyzing 28 frontier large models, we thoroughly demonstrate the detailed performance of current reasoning models across various reasoning questions and provide valuable insights towards constructing more advanced multi-step reasoning models.
@@ F Question Exemplar
G Manual Filtering Details
@@
G.1 Annotation Guideline
G.1.1 Confidentiality Statement
@@ G.2 Annotation Interface
H Human Annotation Details
@@
H.1 Annotation Guideline
H.1.1 Confidentiality Statement
@@ H.1.3 Annotation Standard
H.1.4 Quality Assessment
We adopt a full-scale quality assessment strategy, and the unqualified annotations should be modified until they meet the qualified criteria. We detail the quality assessment process in Section I.
@@ H.2 Annotation Interface
H.3 Details of Video Summary Generation
@@
H.4 Prompts GPT pseudo annotations
@@
I Human Validation Details
I.1 Validation Guideline
I.1.1 Confidentiality Statement
@@ I.1.2 Task Overview
I.2 Configurations of Evaluated Models
We provide the detailed configuration of evaluated VLM, including the model version and input visual frames, which is illustrated in Figure I.2.
@@ I.2 Configurations of Evaluated Models
[…]
@@
[…]
LongVA-7B-DPO 2024-6 LongVA-7B-DPO [88] 128
@@ I.3 Validation Interface
J Evaluation Setup
J.1 Prompts for Model Inference
@@
J.2 Prompts for Open-ended Evaluation
@@
K Ethical Discussion
K.1 Data Safety Discussion
@@
K.3 Potential Social Impact
We argue that VRBench does not contain elements that could pose a potential societal negative impact. We have implemented the aforementioned measures in Section K.1 to prevent the inclusion of harmful information, and we will make every effort to revise the dataset if any content is found to produce unforeseen negative impacts. On the contrary, VRBench might draw the AI research’s focus more to the area of non-Chinese and non-English communities, which we believe is currently being overlooked by published datasets. We consider this to be a potential socially beneficial aspect.
K.4 LLM Usage Claim

### v2

@@ 5 Conclusion
This paper introduces VRBench, a comprehensive long-narrative video benchmark for evaluating multi-step reasoning. Through manually filtered narrative videos, high-quality stepwise annotations, and a multi-phase evaluation pipeline, VRBench distinguishes itself from other existing reasoning benchmarks and shows the robust capability to evaluate LLMs and VLMs both from the process and outcome perspectives. By assessing and analyzing 31 frontier large models, we thoroughly demonstrate the detailed performance of current reasoning models across various reasoning questions and provide valuable insights towards constructing more advanced multi-step reasoning models.
6 Acknowledgement.
This work is supported by the National Key R&\&&D Program of China (No. 2022ZD0160101) and Jiangsu Frontier Technology Research and Development Program (No. BF2024076),
@@ G Question Exemplar
H Manual Filtering Details
@@
H.1 Annotation Guideline
H.1.1 Confidentiality Statement
@@ H.2 Annotation Interface
I Human Annotation Details
@@
I.1 Annotation Guideline
I.1.1 Confidentiality Statement
@@ I.1.3 Annotation Standard
I.1.4 Quality Assessment
We adopt a full-scale quality assessment strategy, and the unqualified annotations should be modified until they meet the qualified criteria. We detail the quality assessment process in Section J.
@@ I.2 Annotation Interface
I.3 Details of Video Summary Generation
@@
I.4 Prompts GPT pseudo annotations
@@
J Human Validation Details
J.1 Validation Guideline
J.1.1 Confidentiality Statement
@@
J.2 Configurations of Evaluated Models
We provide the detailed configuration of evaluated VLM, including the model version and input visual frames, which is illustrated in Figure A3.
@@
Google Gemini 2.0 Pro [61] 2025-2 Gemini 2.0 Pro 0.5fps API
Anthropic Claude-3.7-Sonnet [59] 2025-2 Claude-3.7-Sonnet-No-extended-thinking 20 API
@@
Microsoft Phi-3.5-Vision [1] 2024-7 Phi-3.5-vision-instruct 8 HF
Rhymes Aria [30] 2024-11 Aria-Chat 8 HF
H2O H2OVL Mississippi-2B [17] 2024-10 h2ovl-mississippi-2b 4 HF
DeepSeek DeepSeek-VL2 [79] 2024-12 deepseek-vl2 4 HF
lmms-lab LongVA-7B [91] 2024-6 longVA-7B 128 HF
LongVA-7B-DPO [91] 2024-6 LongVA-7B-DPO 128
Tencent ARC-Hunyuan-Video-7B [archunyuan] 2025-7 ARC-Hunyuan-Video-7B 150 HF
Kwai Keye-VL-8B-Preview [63] 2025-6 Keye-VL-8B-Preview 128 HF
Moonshot Kimi-VL-A3B-Thinking [team2025kimi] 2025-7 Kimi-VL-A3B-Thinking-2506 256 vLLM
Xiaomi MiMo-VL-7B-RL [60] 2025-5 MiMo-VL-7B-RL 512 vLLM
@@
K Evaluation Setup
K.1 Prompts for Model Inference
@@
K.2 Prompts for Open-ended Evaluation
@@
L Ethical Discussion
L.1 Data Safety Discussion
@@
L.3 Potential Social Impact
We argue that VRBench does not contain elements that could pose a potential societal negative impact. We have implemented the aforementioned measures in Section L.1 to prevent the inclusion of harmful information, and we will make every effort to revise the dataset if any content is found to produce unforeseen negative impacts. On the contrary, VRBench might draw the AI research’s focus more to the area of non-Chinese and non-English communities, which we believe is currently being overlooked by published datasets. We consider this to be a potential socially beneficial aspect.
L.4 LLM Usage Claim